OpenStudy (anonymous):

conider the graph defined by y=x-2/x-3 (a) Use the definition of the derivative to find the slope of the tangent line to the graph at the point (4, 2). slope =

ganeshie8 (ganeshie8):

use below to find the slope of tangent to y = f(x), at x = a : slope = \(\large \lim \limits_{x\to a}~ \dfrac{f(x) -f(a)}{x-a}\)

ganeshie8 (ganeshie8):

since you want the slope of tangent line at point (4, 2), start by finding f(4)

ganeshie8 (ganeshie8):

plugin x = 4 in the given function

OpenStudy (anonymous):

f(4)=2

OpenStudy (anonymous):

whats f(a)-?

ganeshie8 (ganeshie8):

good, here \(a = 4\)

ganeshie8 (ganeshie8):

plug that value in and simplify

ganeshie8 (ganeshie8):

slope = \(\large \lim \limits_{x\to 4}~ \dfrac{f(x) -f(4)}{x-4}\) = \(\large \lim \limits_{x\to 4}~ \frac{\dfrac{x-2}{x-3} -2}{x-4}\)

ganeshie8 (ganeshie8):

see if u can simplify^

OpenStudy (anonymous):

-x^2 -x +20

ganeshie8 (ganeshie8):

no, try again

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x\to 4}~ \frac{\dfrac{x-2}{x-3} -2}{x-4} \) \(\large \lim \limits_{x\to 4}~ \frac{\dfrac{x-2-2x+6}{x-3} }{x-4} \) \(\large \lim \limits_{x\to 4}~ \dfrac{-x+4}{(x-3)(x-4)} \) \(\large \lim \limits_{x\to 4}~ \dfrac{-1}{x-3} \) \(\large \ \dfrac{-1}{4-3} \) \(\large -1\)

ganeshie8 (ganeshie8):

see if above makes sense...

OpenStudy (anonymous):

oh yes, sorry i messed up the top part

OpenStudy (anonymous):

so the slope is -1?

ganeshie8 (ganeshie8):

yes !